OpenStudy (anonymous):

What is the inverse of the given relation? Y=7x^2-3

OpenStudy (friendzone):

y=11

OpenStudy (anonymous):

Actually no sorry are you in algebra one or two?

OpenStudy (friendzone):

wait is 7 mulitplying

OpenStudy (anonymous):

I have no clue they Gave me a whole bunch of radicals and exclamations in the answers or I would put them up

OpenStudy (friendzone):

3=14x

OpenStudy (friendzone):

solve y add 3 divide by 14

OpenStudy (anonymous):

A. Y=+ or - sqrt of x+3/7 stuff like this

OpenStudy (anonymous):

y+3=7x^2 x^2 = (y+3)/7 x= sqrt((y+3)/7) so we say : y=sqrt((x+3)/7)

OpenStudy (anonymous):

this is the inverse.. (we solved for x)

OpenStudy (anonymous):

y=+/-sqrt((x+3)/7)

OpenStudy (anonymous):

just interchange y and x then solve for y!! y=sqrt((x+3)/7)

OpenStudy (sheraz12345):

\[y=7x ^{2}-3\] \[y-3=7x ^{2}\] \[\sqrt{\frac{ y-3 }{ 7 }}=x ^{2}\] So answer will be \[\pm \sqrt{\frac{ x-3 }{ 7 }}=y\]